The sum of a number and 2 less than 1/2 the number is 30. Find the number.

Answer by macston(5194)

.
x=the number
(0.5x-2) = 2 less than 1/2 the number
.
x+(0.5x-2)=30
1.5x-2=30
1.5x=32
x=21.33
.
ANSWER: The number is 21.33
.